OpenStudy (saifoo.khan):

Completing square with steps!!! PLEASE. -0.035x^2 + 1.4x+6 = 0

OpenStudy (amistre64):

the step is simply; half the coeff of the x term and square it

OpenStudy (amistre64):

1.4/2 = .7, ^2 = .49

OpenStudy (saifoo.khan):

then?

OpenStudy (saifoo.khan):

others said me to divide the whole equation by -0.035

OpenStudy (amistre64):

then add and subtract that amount into the equation

OpenStudy (amistre64):

..... i missed that dint i

OpenStudy (amistre64):

yeah, you have to factor off the leading coeff first ...

OpenStudy (saifoo.khan):

x^2 -40x -171.428571 = 0

OpenStudy (amistre64):

-0.035x^2 + 1.4x+6 = 0 -0.035(x^2 -1.4x/.035) +6 = 0 -0.035(x^2 -1.4x/.035 +(1.4/.035(2))^2)-(1.4^2/.035(4)) +6 = 0 maybe

OpenStudy (amistre64):

its not that it difficult, its just messy

OpenStudy (saifoo.khan):

I just want to put it in completing square form, the answer is: -(0.035) (x - 20)² + 20

OpenStudy (saifoo.khan):

Sat tried to help, but his answer wasn't correct.

OpenStudy (amistre64):

let me see if i can mutilate it from the start :)

OpenStudy (saifoo.khan):

Thank you a ton!

OpenStudy (amistre64):

-0.035x^2 + 1.4x+6 = 0 -0.035x^2 + 1.4x = -6 x^2 + 1.4x/-0.035 = -6/-0.035 x^2 - 1.4x/.035 = 6/.035 x^2 - 1.4x/.035 +(1.4/.035(2))^2 = 6/.035 +(1.4/.035(2))^2 (x - 1.4/.035(2))^2 = 6/.035 +(1.4/.035(2))^2 (x - .7/.035)^2 = (6(.035)+.49)/.035^2 (x - .7/.035)^2 = (.21+.49)/.035^2 (x - .7/.035)^2 = .7/.035^2 (x - 20)^2 = .7/.035^2 (-.035)(x - 20)^2 = .7(-.035)/.035^2 (-.035)(x - 20)^2 = -.7/.035 (-.035)(x - 20)^2 = -20 (-.035)(x - 20)^2 +20 = 0

OpenStudy (saifoo.khan):

im trying to read. You are awesome btw. ;)

OpenStudy (amistre64):

i am arent i ... lol